A= 1/2h (b1+b2 for b1 Please help,
solve the problem for the indicated letter

Found 2 solutions by solver91311, Alan3354:
Answer by solver91311(24713)





Multiply by 2:



Multiply by



Add to both sides:




John


Answer by Alan3354(69443)

A= 1/2h (b1+b2 for b1
------------
A right bracket is missing, I think it goes here:
A= 1/2h (b1+b2)
2A = h*(b1 + b2)
2A/h = b1 + b2
b1 = (2A/h) - b2